OpenStudy (anonymous):

Find all solutions of the equation on the interval 0 degrees < x < 90 degrees. sin (x + 30degrees) - root3 sin x

OpenStudy (anonymous):

\[\sin (x+30degrees) = \sqrt{3} \sin x\]